# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v1

@@ Appendix A Elastic and inelastic scattering in a tight-binding, isolated one-dimensional chain
Multi-phonon and hybrid CM-phonon processes
When single-phonon processes are kinematically suppressed, higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange are allowed. These channels relax the strict resonance condition Eq. (47) at the cost of reduced matrix elements.
Kinematic summary and temperature dependence
  • Elastic channels: exact backscattering $k\rightarrow-k$ without phonon excitation occurs via CM recoil and is enabled by translational invariance.
  • Inelastic channels: internal phonon emission and absorption become increasingly important as temperature increases and as phonon spectra permit energy matching.
  • Finite size: the discreteness of the phonon spectrum in a finite chain strongly constrains available inelastic channels and can enhance the relative importance of CM-mediated elastic processes.
@@
  • In wavepacket simulations, elastic backscattering events are identified by momentum reversal without change in internal phonon occupation, accompanied by a compensating shift in CM momentum.
  • Both bulk and boundary reflections obey the same momentum conservation law and are distinguished only by their spatial location.
  • The CM zero mode plays a central dynamical role and cannot be neglected without qualitatively altering the scattering physics.

### v2

@@ A Golden-Rule rate has the form
Multi-phonon and hybrid CM–phonon processes
Higher-order multi-phonon processes and hybrid processes involving both CM recoil and internal phonon exchange relax the strict resonance condition (47) at the cost of reduced matrix elements.
Kinematic summary
  • Phonon-diagonal channel: $k\to-k$ with unchanged internal state is allowed by translational invariance and proceeds via CM recoil; it is strictly elastic only on resonance (47) and effectively elastic for large $N$.
  • Inelastic channels: internal phonon emission/absorption become important when the phonon spectrum permits energy matching.
  • Finite size: discrete internal spectra constrain inelastic channels and can enhance the relative importance of CM-mediated phonon-diagonal scattering.
@@
  • In wavepacket simulations, phonon-diagonal backscattering is identified by momentum reversal with unchanged internal state and a compensating shift of $P_{R}$ by $2\hbar k$.
  • Bulk and boundary reflections differ only in spatial location; both satisfy the same momentum conservation enforced by the CM zero mode.
  • Neglecting the CM zero mode (implicit clamping) qualitatively changes the available scattering channels.
Appendix C Whither Umklapp?
The traditional role of Umklapp for finite resistivity in an infinite, perfect crystal arises from freezing the background and projecting out its pseudomomentum. Once the background is treated as dynamical, with an explicit translational degree of freedom, momentum relaxation can occur through elastic pseudomomentum exchange between electron and lattice. Phonons then represent one possible basis for describing background dynamics rather than the exclusive channel for momentum transfer.
This perspective clarifies long-standing ambiguities surrounding Umklapp and explains why momentum relaxation is observed experimentally even in regimes where Umklapp processes are expected to be suppressed. It also underscores that the association of Umklapp with dissipation and entropy production is not intrinsic, but follows from additional assumptions about the background.
The expectation that Umklapp processes are necessary for finite resistivity in a perfect crystal has deep roots in transport theory. Peierls first distinguished between normal and Umklapp processes in his analysis of phonon thermal conductivity, showing that only the latter can relax total crystal momentum in a closed system. This reasoning was later extended to electrons by Bloch, Peierls, Ziman, and others, and appears explicitly in standard texts (e.g. Peierls, Quantum Theory of Solids; Ziman, Electrons and Phonons). If the Fermi surface lies well inside the Brillouin zone, or at low temperature where large-wavevector phonons are thermally suppressed, Umklapp processes are expected to be exponentially rare. Under these conditions conventional theory predicts extremely long momentum lifetimes in a clean crystal.
The present formulation revisits this conclusion by restoring the dynamical translational degree of freedom of the background. Momentum exchange between subsystems does not require phonon creation, and the strict identification of Umklapp with momentum relaxation relies on additional assumptions about background clamping or environmental coupling.
Umklapp processes are best understood as a classification of crystal-momentum conservation rather than as the fundamental cause of momentum relaxation. Their elevated role in conventional transport theory stems from treating the lattice background as inert. Recognizing the background as a dynamical object carrying its own pseudomomentum restores a more symmetric and physically complete description in which elastic momentum exchange is allowed and Umklapp is no longer indispensable.